Consider the independent investment projects in the table below. Compute the project worth of each project at the end of sixProject Cash Flows oo AWN-Os - $800 150 150 350 - 200 500 400 - $500 150 150 150 $200 - 40 -60 - 140 100 100 ||Single Payment Compound Present Amount Worth Factor Factor (F/P, i, N (P/F, I, N 1.1000 0.9091 1.2100 0.8264 1.3310 0.7513 1.Əo o voo AwN+z. Single Payment Compound Present Amount Worth Factor Factor (F/P, i, N) (P/F, I, N) 1.1500 0.8696 1.3225 0.756
Answer #1

Present worth of project A = - $ 800 + $ 150 \times (P/F, 10%, 1 year) + $ 150 \times (P/F, 10%, 2 year) + $ 350 \times (P/F, 10%, 3 year) - $ 200 \times (P/F, 15%, 4 year) + $ 500 \times (P/F, 15%, 5 year) + $ 400 \times (P/F, 15%, 6 year)

Present worth of project A = - $ 800 + $ 150 \times 0.9091 + $ 150 \times 0.8264 + $ 350 \times 0.7513 - $ 200 \times 0.5718 + $ 500 \times 0.4972 + $ 400 \times 0.4323

Project A worth at the end of 6 years will be $ 30.44

----------------------------------------------------------------------------------

Present worth of project B = - $ 500 + $ 150 \times ( P/A, 10%, 3 years) + $ 100 \times (P/F, 15%, 4 years) + $ 100  \times ( P/F, 15%, 5 years)

Present worth of project B = - $ 500 + $ 150 \times 2.4869 + $ 100 \times 0.5718 + $ 100 \times 0.4972

Project B worth at the end of 5 years will be $ - 20.07

--------------------------------------------------------------------------------

Present worth of project C = $ 200 - $ 40 \times ( P/F, 10%, 1 year) - $ 60 \times (P/F, 10%, 2 years) - $ 140 \times (P/F, 10%, 3 years)

Present worth of project C = $ 200 - $ 40 \times 0.9091 - $ 60 \times 0.8264 - $ 140 \times 0.7513

Project C worth at the end of 3 years will be $ 8.87